OpenStudy (anonymous):

A circle has its center at (−2, −3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y − 3)2 = 9 (x − 2)2 + (y − 3)2 = 3 (x + 2)2 + (y + 3)2 = 9 (x + 2)2 + (y + 3)2 = 3

OpenStudy (anonymous):

so the formula will be (x-x1) + (y-y1) = r^2

OpenStudy (welshfella):

general form is (x - a)^2 + (y - b)^2 = r^2 where (a,b////0 is the centre and r = radius

OpenStudy (anonymous):

so first of all, the -2 would be your x1

OpenStudy (welshfella):

* (a,b) is the centre

OpenStudy (anonymous):

making the first part (x+2)

OpenStudy (anonymous):

so what would you think the second part would be?

OpenStudy (anonymous):

oh i screwed up, the (x+2) would be squared too

OpenStudy (welshfella):

( y - (-3))^2 = ?

OpenStudy (welshfella):

what does a double negative give you?

OpenStudy (anonymous):

and you have a radius of 3, so the number on the right side would be three squared

OpenStudy (anonymous):

so what does that make?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

close

OpenStudy (anonymous):

remember that your y coordinate is at -3

OpenStudy (anonymous):

and the formula suggests subtracting your y coordinate

OpenStudy (anonymous):

so you're subtracting a negative

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np